OpenStudy (anonymous):

Joe climbs a 5 foot ladder to the top of a slide. Find the length of the slide, to the nearest hundredth, if the distance from the base of the ladder to the bottom of the slide is 9 feet. Note: The ladder forms a right angle with the ground. is it 7.48ft, 8.43ft , 10.30ft or 11.45ft

OpenStudy (amistre64):

what type of engneer designs a ladder that is 90degrees to the ground; thats just a lawsuit waiting to happen ..

OpenStudy (amistre64):

d^2 + h^2 = s^2 9^2 + 5^2 = s^2 81 + 25 = s^2 ; you can work it from there right?

OpenStudy (anonymous):

yes